OpenStudy (anonymous):

Which formula? O.o @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you'll use A = P*(1+r)^t where A = final value after t years P = initial value r = growth/decay rate (positive for growth, negative for decay) t = time in years

OpenStudy (anonymous):

209,350*(1+0.045)^15= A ?

OpenStudy (anonymous):

If so, A= 405151.379 or .38

jimthompson5910 (jim_thompson5910):

getting the same and yes you round to the nearest penny